1. Which gas is most dense at 1.00 atm and 23·C? CO, (2 points) .N20, or Cla. Show work and explain. 2. Consider the following reaction: (5 points) CaC2(s) + 2 H2O(1) ? Ca(OH)2(s) + C2H2(g) a) How many grams of CaC2 are needed to generate 112 L of C2H2 gas if the pressure of CaHz is 199.6 kPa at 60 c? b) Using the above equation, how many Liter of CzHz gas are produced if 966 mL of water are used at 612 torr and 85°C?
Answer #1

1)

1)

Lets derive the equation to be used
use:
p*V=n*R*T
p*V=(mass/molar mass)*R*T
p*molar mass=(mass/V)*R*T
p*molar mass=density*R*T

So, the one with greater molar mass will be most dense


Molar mass of CO2,
MM = 1*MM(C) + 2*MM(O)
= 1*12.01 + 2*16.0
= 44.01 g/mol

Molar mass of N2O,
MM = 2*MM(N) + 1*MM(O)
= 2*14.01 + 1*16.0
= 44.02 g/mol

Molar mass of Cl2 = 70.9 g/mol

Clearly, molar mass is highest for Cl2

So, Cl2 is most dense

Only 1 question at a time please